How many moles are in 2.3 g of NaCl?
1
Expert's answer
2019-12-22T02:38:27-0500

Molar mass( NaCl)=23+35.5=58.5g/mol

number of moles(NaCl)=2.3g/58.5g/mol=0.039 mol